We observe the following market data for the U.S. and Switzerland. The Fisher relationship holds.
• annual U.S. inflation = ??? • one-year U.S. interest rate = 3.95%
• annual Swiss inflation = 0.12% • one-year Swiss interest rate = 0.80%
The spot SFr/$ is 1.2306. What is the annual U.S. inflation? What is the expected spot SFr/$ in one year according to uncovered interest parity?
Answer #1

Using PPP & IFE,

[(1+IUS)/(1+IS)] - 1 = [(1+iUS)/(1+iS)] - 1 or

(1+IUS)/(1+IS) = (1+iUS)/(1+iS) where

IUS = US inflation; Is = Swiss inflation =0.12%; iUS = US interest rate = 3.95%; iS = Swiss interest rate = 0.80%

1 + Ius = (1+3.95%)*(1+0.12%)/(1+0.80%) = 1.0325

Ius = 1.0325 -1 = 0.0325 or 3.25%

As per uncovered interest rate parity,

F(SFr/$) = S(SFr/$)*(1+ iS)/(1+iUS)

= 1.2306*(1+0.80%)/(1+3.95%) = SFr 1.1933/$